July 18, 2019 Report - July 8 - Aug 2
Director - Todd Satogata
Previous two weeks (July 8 - July 19) - ADMIN: Travel, FY account planning - ADMIN: DOE S&T review (Jul 11) and prep - ADMIN: Max Bruker arrival planning - ADMIN: Safety all-hands meeting (Jul 19) - JLEIC: pCDR development, EICUG talk prep, coordination meetings - JLEIC: parameter tables, lattices to baseline - ODU: REU, graduate student mentoring - ODU: Courtney Washington meeting (Jul 19) - JACOW: ICALEPCS'19 invited talk coordination Next two weeks (July 22 - August 2) - ADMIN: Travel, FY account planning - ADMIN: Academic writing workshop planning (Jul 24) - ODU: REU student mentoring, ODU presentation (Jul 26)/poster/paper prep - JLEIC: pCDR development, Aug 2 submission prep, coordination meetings - JLEIC: parameter tables coordination
Previous two weeks (July 8 - July 19) - Machine support - Expand USPAS lecture on Radiation Damping - Revisions to Operations Training, MOODLE - Final review of IOP submissions for the IPAC'19 Light Peer Review Next two weeks (July 22 - Aug 2) - Vacation - PRAB paper review
Previous two weeks (July 8 - July 19) - Participate the weekly ALS-U impedance meeting - Interesting discussions with Karl Bane and Bob Warnock at the meeting - Bob Warnock is modifying his code on microwave instability so that he can send to me to run for JLEIC case - Prepare for JLEIC impedance meeting on 6/17: - Talking to Tim Michalski to prepare presentation on vision and status of the engineering design - Talking to Rimmer to prepare presentation on vision and status of the feedback system - Uneven bunch spectrum - Write up technote for my analysis, - start to look at its impact on the coupled-bunch instability growth rate - Plasma cascade in the BNL Coherent-electron-cooling - Read reference papers on, trying to catch up on their current status - Remotely participated one CeC operation meeting on their upcoming experiment - Assisted Walter Wittmer on R&D list - Work with Bob Rimmer to prepare documents Walter requested for the R&D list, in particular, I'm the P.I. of block 6 (impedance and instabilities) in the list - Assisted with CBETA on CSR induced microbunching instability - Georg's graduate student, William Lou, is following our suggestion to do convergence test of the microbunching instability. He had questions for me and my former student Cheng-Ying Tsai. And we are helping him on the problems. Next two weeks (July 22 - Aug 2) - Impedance studies - Organize the JLEIC impedance meeting on 7/16 - Participate ALS-U impedance meeting and discussions - Continue to understand impedance analysis - Help to coordinate impedance budget study at JLAB - Coupled bunch instability - Continue with technote writing - Continue with analysis of coupled bunch instability and the study of the impact of uneven bunch distribution on growth rate - CeC project - Continue to understand the underline physics - Continue to participate their weekly meeting - CSR in CBETA - Continue to give support/help on their CSR studies
Previous two weeks (July 8 - July 19) - Worked on tuning the LEB - Have been performing some duties as optics on call - Have been working on my proposal for the DESIREE experiment. Next two weeks (July 22 - Aug 2) - I will finish and submit the proposal to the University of Stockholm - Finish up my tenure as Optics on Call - Finish up my redesign work on the two boosters
Previous two weeks (July 8 - July 19) - Final LDRD slide prep and presentation - BNL CeC telecon presentation on greenfield ERL design - LERF Ops (single day for isotope run) - Work on JSA initiative proposal - Bagels with Andrei - Interview panel for Machine Learning Programmer Analyst posting - Leading ML Lunch Series - Lay out 476 MHz version of linac for JLEIC ERL cooler Next two weeks (July 8 - July 19) - Propagate distribution from Fay through (completed) parts of JLEIC ERL cooler - BNL CeC analysis of greenfield front end design - Work on RF classifier pipeline
Accelerator R & D - Yuhong Zhang
Previous two weeks (June 24 - July 5) - Developed estimated luminosity curves for 140 GeV JLEIC with or without high energy cooling - Worked on a concept for 140 GeV JLEIC - Read chapter 3 of pre-CDR-100 and made corrections - Worked on three talk slides. Next two weeks (July 8 - July 19) - Continue work from previous weeks
Previous two weeks (July 8 - July 19) - Relocation - Orientation and training - Read some JLEIC-related material Next two weeks (July 22 - Aug 2) - Shake some more hands - Familiarize myself further with JLEIC
Previous two weeks (July 8 - July 19) - Working on updating the IR design in the electron collider ring - Working on the transparent spin project - Working on the storage ring cooler project - Preparing tutorial material for ZGOUBI workshorp - Miscellaneous: attended the Innovative HE Cooling Project Status meeting, attended the Storage Ring Cooler meeting, attended JLEIC R&D meeting, attended ZGOUBI workshop committee meetings, attended JLEIC general meeting, attended IP magnet design meeting, attended teleconference on transparent spin test meeting, attended Packing factor/drift limit program meeting Next two weeks (July 8 - July 19) - Work on updating the IR design in the electron collider ring - Work on the transparent spin project - Work on the storage ring cooler project - Prepare tutorial material for ZGOUBI workshop
Previous two weeks (July 8 - July 19) - Working on elegant simulation using real parameters for RF cavities instead of using matrices. - Introducing real lattices in elegant simulation instead of using matrices. - Understanding the theory on damping time and energy spread in two energy storage ring ( Dr. Krafft's tech note). - Making more progress in stability condition in two energy storage ring. - Enjoyed couple of days in New York. Found that CASA, Jefferson Lab is the best place for me to enjoy. Next two weeks (July 22 - Aug 2) - Come up with the final longitudinal stability in two energy storage ring with real lattices and real cavities. - Introducing the damping element in simulation, calculate the damped emittance. - Understand the concept of multiple energy storage ring (Dr. Krafft's new idea). - Start to think about NAPAC2019.
Previous two weeks (July 8 - July 19) - Transparent spin experiment: spin tracking simulations, error analysis, teleconference with BNL collaborators - Study of the 400 GeV ion collider ring upgrade: IR redesign, crab crossing estimates and polarization control scheme - Support of Walter's work Next two weeks (July 22 - Aug 2) - Preparation for the CFNS summer school - Preparation of DoE quarterly reports - Updating the pre-CDR - Spin and beam dynamics simulations
Previous two weeks (July 8 - July 19) - Parameterization of the probability distribution for exit energies and angles for 13.1 MeV/u lead ions passing through carbon foils as a function of foil thickness and incident beam energy. Bilinear interpolation of the probability distribution to account for variance in the foil thickness and energy spread in the incident beam. - Mapping of the distribution for ions before the stripping foil to a distribution accounting for the energy straggle and growth of angular spread due to multiple scattering effects in the stripping foil. - Initial simulations in TRACK of stripping effects and transport of lead ion beams containing multiple charge states; previous simulation included beam only stripped to a single charge state, and did not include energy or angular spread effects in the stripped beam. - Some preliminary upstream stripping chicane elements implemented in MADX for initial beam optics matching. Next two weeks (July 22 - Aug 2) - Continued work on the distribution mapping for pre-strip to post-strip lead ion distributions, and stripping chicane optics design and matching to the stripped beam distribution. - Continued work with PyORBIT including: troubleshooting the DTL linac implementation and lack of beam stability, and optimization of the superconducting portion of the linac for better matching with TRACK. Also plan to discuss with ANL the possibility of converting the RF cavity field map files from binary Fortran files to something that can be read into PyORBIT. - Also the possibility of implementing the stripping chicane in PyORBIT for more flexibility in handling modified beam distributions - TRACK does not appear able to do this.
Computational Physics - Yves Roblin
Previous two weeks (July 8 - July 19) - CEBAF: - Bteam, optics on call, machine setup - Technote writing on multiscanner emittance and beam properties determination in hall A. - JLEIC: - Quarterly report on beam-beam tools - Investigating non linear RF transport in BB3D Next two weeks (July 22 - Aug 2) - More CEBAF optics support - Simulation of bbeam kinematics for pCDR
Previous two weeks (July 8 - July 19) - Scaled SuperKEKB FFQ multipole data - Compared results obtained by different methods used to compensate chromaticity and found a larger DA than previous result. - Dynamic aperture with multipoles on one upstream FFQ using the SuperKEKB data - Maximize bare lattice DA using tune (in progress) [tune: A program to apply the MINUIT minimization engine to an arbitrary script.] Next two weeks (July 22 - Aug 2) - Maximize bare lattice DA using tune (in progress) - Continue dynamic aperture studies with multipoles.
Previous two weeks (July 8 - July 19) - Spin Transparency Mode - For Spin Tracking (±𝟎° for two snakes, no IP bumps, 𝒔𝒚𝒏𝒄𝒉𝒓𝒐𝒏𝒐𝒖𝒔 𝒑𝒉𝒂𝒔𝒆=𝝅), the effects of adjusting the reference closed orbit length have been teased and checked. Involving the configuration for accelerations, the different configurations of Zgoubi parameter, “CAVITE”, have been tested. - For Spin Tracking (±𝟎° for two snakes, no IP bumps, 𝒔𝒚𝒏𝒄𝒉𝒓𝒐𝒏𝒐𝒖𝒔 𝒑𝒉𝒂𝒔𝒆=0) using different RHIC’s configurations, we have carried out the precise calibration on ALE angle and B1-field for PHIC’s configurations based on Vasiliy’s analytic solution. Now the revised RHIC’s configuration will yield the spin transparent. - For Spin Tracking (±𝟎° for two snakes, with chromaticity’ compensation), Y-shift, Z-shift and the skew angle changes have been introduced in RHIC’s configurations. The effect of the above shifts has been studied both in the statistic results and Zgoubi simulation results. - For Spin Tracking (±𝟎° for two snakes, with chromaticity’ compensation), the closed orbit was studied. - The research progress was reported on the teleconference on transparent spin meeting (July 18, 2019) - Program codes for Zgoubi - As a supplementary of Zgoubi, we designed and completed the following Python based program codes: - ALE_and_B1_Calibration.py - ChangeREF_DY_DZ_for_Multipole_and_SkewAngle_for_Bend.py - Closed_orbit_sigma_ALE_B1_and_DY_DZ.py - Closed_orbit_sigma_ALE_B1_and_DY_DZ_and_Dphi.py Next two weeks (July 22 - Aug 2) - Keep working on the project of “Theoretical and Experimental Study of the Spin Transparency Mode in an EIC”.
Previous two weeks (July 8 - July 19) - Compared the luminosity output for all proposed energy values for JLEIC using BB3D and GHOST and calculated their relative differences. - Performed convergence studies to find the best slice number for hourglass simulation. - Calculated the hourglass reduction using GHOST for CDR44.7 GeV Next two weeks (July 22 - Aug 2) - Perform gear changing simulations required for the submitted abstract of NAPAC2019 - Study the beam-beam dynamics with gear-changing
Previous two weeks (June 24 - July 5) - Cooling code development - A model for sweeping effect has been added to JSPEC. We first calculated the maximum friction for the given ion beam, electron beam and electron cooler. Then we calculate the proper acceleration of the electron beam for the sweeping process and the time needed for it. Once the motion of the electron beam is is determined, the sweeping process is simulated in n steps, where n is a number defined by the user. In each step, the friction force on all the ions are calculated with the increased relative velocity due to the motion of the electron beam and the kicks are applied on all ions. In the test run, the sweeping effect has been successfully observed. But there are some issues left in the test: (1) cooling is artificially strengthened for demonstration; (2) transverse temperature of the ion beam is assumed to be zero; (3) synchrotron motion is ignore; (4) IBS effect is ignored. Will include these effect one by one later in the simulations. Coding of the IBS formula by Kubo and Oide is finished, but the result is wrong. Has not figure out it is a bug in the code or any mistakes in the methodology that I implied. - Beam-beam effect - Revised our C++ symplectic tracking code and shipped it to Yun Luo. - Others - Prepared the talk for my LDRD application and present to the LDRD review panel. Next two weeks (July 8 - July 19) - Add symplectic tracking to BB3D and verify the codes together with Yun Luo. - Continue sweeping effect simulation. - Abstract for COOL’19 conference.
Diagnostic Development - Kevin Jordan
Previous two weeks (June 24 - July 5) - Next two weeks (July 8 - July 19)
Previous two weeks (July 8 - July 19) - HPWS: The linear motor controller is undersized for the HWPS when the stage in at 45 degrees. This angle is the nominal operation angle. A collaborator from Radiabeam was here for a couple of day to try tuning the drive for better performance. This didn’t work and a larger driver was ordered. The HPWS was moved to the optical table to allow operations in the horizontal position. This allowed us to test the EPICS controls and to obtain scan data for a student summer project. - Isotope: Helped with aligning and testing the diagnostics. Worked with operations during the two day run. At the end we were able to thread beam to the 1G dump. - FAST: Designed the tungsten slit plate, new fork and gimbal plate. The gimbal plate will allow the fork to be set normal to the beam. The new fork and gimbal plate have been made and are ready for cleaning. Order titanium bolts to replace the silver plated bolts in the LDRDS. Next two weeks (July 22 - Aug 2) - HPWS: Use the new dynamic tuning SW to improve performance before the larger controller arrives. Test the incremental encoder with a long cable. - Isotope: Complete the alignment and installation of the shield wall. Support the upcoming review. - FAST: Supervise the disassembly, cleaning and partial assembly of the LDRDS. - Halo Monitor: Align and install a fiber coupler and camera assemble to be able to obtain spectrum from the electron/BNNT interaction.
Previous two weeks (July 8 - July 19) - Spent essentially all of my time on real-time support of the accelerator tuning and improvements to the Hall B setup process (see below). Plus a few days of vacation leave. - The envelope-based matching scheme I proposed to improve accelerator tuning (prospectively supported by the Arc 7A SLM complement, of which the 7A08 sync. light monitor is the first) is the motivation of the process I initiated with operators during tuning of the accelerator this round. It became clear that with the Spreader matching quadrupoles "on design," empirical adjustment of the defined envelope tuning knobs of the previous Recombiner to make the viewer spots of the following recirculation arc most alike was a protocol which matched the beam across the linac in a "best practically achievable" sense. This provides minimally perturbed beam optics across the linac and through the Spreader, minimizing time in execution to make the qsUtility scans and re-match through the Spreader. Boron Nitride NanoTube material, as recently installed at viewer ITV7A09, may also provide a sufficiently well-resolved image of the beam (although invasive in image acquisition) to serve this purpose. This may be a feasible alternative to distributing SLMs throughout the arcs, sidestepping the characteristic blooming limitation of the original ChromOx viewer material which has prevented application of the original (late '80s) design intent of using the viewers in this way. Next two weeks (July 22 - Aug 2) - Scheduled leave for the following two-week calendar period
LERF - Steve Benson
Previous two weeks (July 8 - July 19) - Worked on finalizing documentation for Isotope ERR. - Worked on quarterly summaries for FOA contract and budget for Isotope quarterly. - Assisted in running the LERF beam to the 1G dump to check out the Beamline hardware for isotope production. - Set up two August Accelerator Seminars. - Reviewed RSI paper. Next two weeks (July 22 - Aug 2) - Get LDRD sent to Fermilab. - Finalize documentation for isotope ERR - Complete quarterlies for the FOA contract and the budget for the isotope project. - Continue to work on updating the Pre-CDR-100 chapter 6.
Previous two weeks (July 8 - July 19) - Attended the SuperKEKB Review in Japan July 8-10. A lot of information that is very relevant to JLEIC, particularly in the background and collimation topics. All the talks are available on the web at http://www-kekb.kek.jp/MAC/. - Then a few days vacation in Koyasan - This week in Isotopes, we have got pulsed beam to the 1G dump, still need CW (awaiting the MPS masks) and then getting the beam to our target, the 1X line. We have made considerable progress on the equipment to unscrew the carbon plug in the back of the crucible (Kevin has built a turntable to rotate the crucible in the hot-cell while the remote manipulator is used to hold the plug). Everything is now ready for a trial transfer to VCU. We will pack the crucible into the pig into the shipping container today. - I did a little editing of the pCDR. I also spent some time with Sarah. Next two weeks (July 22 - Aug 2) - We (Kevin, Joe and Keith Welch) will go up to VCU on Monday to test the unpacking of the pig from the shipping container (with VCU RadCon) and then opening the pig in the hot-cell using remote manipulators (with Jamal’s group. The two students from NMT, Geno Santistevan and Robert Bentley, will be here next week. They will be doing thermal transfer measurements and getting up to speed on everything we have been doing. I have asked them to give us a presentation of their work so far. - I will also be doing a lot of pCDR editing. - This is Sarah’s last two weeks, so she might need help finishing her project. The poster and paper are already partially written.